

********************************************************************
UARS WEEKLY STATUS REPORT - 24 SEP 1999
********************************************************************
***********
SOLSTICE
***********
SOLSTICE continues to operate normally, collecting solar data on all
available orbital tracking periods and stellar calibration observations,
whenever possible.
***********
SUSIM
***********
The gridded offset maneuver performed last weekend mistakenly repeated
one set of offsets. The impact is being assessed. Otherwise, SUSIM
gathered its normal solar and occultation data this past
week without apparent problems.
***********
ACRIM
***********
Nothing received.
***********
PEM
***********
The PEM spectrometers continue to be active under PMON control. While
under PMON control, the PEM instrument computed energy deposition from its
particle spectrometers.
The FOT and the PEM team are updating the PEM activation sequence which
effects the PEM/HEPS (> 30 keV particles). Current activation sequence
reduces the activation time by about 1 minute. Additional analization shows
that the activation sequence may be reduced by about another minute. PEM
PMON activation is set for the beginning of power orbit day. The beginning of
power orbit day is defined as when:
array current greater than 5 amps, and
array voltage greater than 53.5 volts, and
any battery current greater than 2.0 amps
The end of power orbit day (PEM deactivation) is defined as when:
time in power orbit day greater than 4 minutes, and
array current less than 9 amps, and
all battery currents less than 0.2 amps
Power orbit day periods vary depending upon beta angle:
about 39-40 minutes at yaw
about 50-51 minutes at max beta
Current activation takes about 7.5 min for full activation. During the
activation sequence, various components of PEM become active at all different
times, beginning with the PEM magnetometer (at about on activation time) and
ending with the HEPS-2 components (7.5 min later).
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
WINDII is next scheduled to operate on October 12-23.
***********
MPG
***********
Routine processing of instrument Daily Activity Plans continues
with no problems. The MPG continues to assist the Project
Scientist with the scheduling of instrument operations.
***********
CDHF
***********
The CDHF processed current day data through level 3 for all active
instruments. Additionally, CDHF personnel reprocessed: (R)= Requested
(C) = Complete, (F)= Failed (L) = Days remaining to reprocess.
13 days CLA (7.0) PROD (L1-L3) (473R-269C-16F=188L)
0 days GPI (4.05) PROD (CLAES L3B) (191R-164C-2F=25L)
0 days WIN (5.11A)TEST (L1-L3) (3R-0C-0F=3L)
The CDHF has PCMB approval for the configuration and test of HALOE V4.82
formal delivery. WINDII 5.11A formal delivery needs PCMB approval for
configuration. Three days are requested in test mode.
CDHF personnel continued to draft procedures for the new Alpha 4000 and
HSM systems. We have begun testing the new Rehost system by running
the Production Program "Attitude Extraction" and performing a Level 0 Ingest.
Once we have these successfully running we shall start on the PI Production
Programs.
***********
FOT
***********
REPORTING PERIOD
This report covers 11 September, 1999 (Orbit 43717, GMT Day 254) through 17
September 1999 (Orbit 43820, GMT Day 260).
SPACECRAFT OPERATION
The observatory is now in Reverse Flight (Northern Hemisphere viewing) and is
performing nominally under two-battery operations. The instrument operational
changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
WINDII 257/2218 43775 Off
The beta angle ranged between the angles of 49.5 degrees to 27.4 degrees for
this report period. The beta angle is now decreasing toward a minimum angle of
-2 degrees on 24 September 1999 (DOY 267).
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 23.8 V end-of-night (EON) load bus voltage. Battery
temperatures are stable with a temperature delta between Battery 2 and Battery
3 of 1.0 to 1.3 degrees C. Additional battery performance data for each
battery on 10 September 1999 and 17 September 1999 are provided in Appendix D.
The clock error ranged from 10.3 to -10.8 msec during this report period.
Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial excursion of
ALERTS or Out-Of-Limits conditions detected by the FOT for this report. A
summary is listed below and the details of ALERT and Out-Of-Limits occurrences
are listed in Appendix B.
% ALERTS = 00
% OUT-OF-LIMITS = 02
The following real-time command sheets were executed for the listed operational
element(s) during this report period (see Appendix A).
% FOT = 04
% WINDII = 01
UARS Anomaly Reports initiated during this report period are listed in Appendix
E and summarized below.
% FOT = 03
Four (4) Anomaly Report remained OPEN at the end of this report period (see
Appendix E).
TRANSPONDER FREQUENCIES:
% Transponder A: 99/242 - 2287.496953
% Transponder B: 99/243 - 2287.499030
GROUND SYSTEM OPERATION
Three (3) operational support problems occurred this period. Anomaly Reports
are included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from mission
start through UTC Day 99/259 show the observatory data loss to be 24 hours, 55
minutes, 17.108 seconds (an increase of 7.680 seconds since the last report
period). This is a 0.0356 percent data loss which equals a 99.9644 percent
data capture for the mission.
The increase of 7.680 seconds occurred over one (1) day:
7 7.680 seconds on Day 257 due to an IP line hit.
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
PEM requested to have the timing in their turn on sequence tightened up in yet
another area. This is expected to be accomplished within the next few weeks.
YEAR 2000 OPERABILITY SUPPORT
End-to-end Y2K testing continued this past week. The FOT received the Brower
Mean Element report from FDF, but are still waiting for a valid 8-week PSAT
file to be used for testing. Without a complete set of 8-week files, the UARS
CMS is unable to perform all of the necessary mission planning and load
building processes required for Y2K end-to-end testing. In the meantime, the
FOT has generated a "test" Y2K ATC load by editing an existing 1999 load, and
performing a universal substitution edit command on the year, changing it from
1999 to 2000. This ATC load was tested on the simulator and worked properly.
MEETINGS
None
FUTURE SIGNIFICANT EVENTS
A meeting to discuss the UARS Safehold configuration and concerns for the MPS
and its batteries is planned for September. This meeting will be a closed
meeting with only the FOT and specialized engineering support attending.
PERSONNEL CHANGES
None.
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 17 September 1999
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ -------
43724 254/1158 99-320 OBC Clk Adj fr 0.00 to 20.6 msec/day
43751 256/0703 99-321 OBC Clk Adj fr 20.6 to 0.00 msec/day
43775 257/2218 99-322 WINDII Power OFF
43794 259/0442 99-323 OBC Clk Adj fr 0.00 to 20.6 msec/day
43814 260/1159 99-324 OBC Clk Adj fr 20.6 to 0.00 msec/day
__________________________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 17 September 1999
MNEMONIC STATE DESCRIPTION ORBITS
--------- ----- ------------ ------
HRPTSPBIMON Y-HI HRDI PTS PB Current Mon 43792,97
HRPSEQBIMON Y-HI HRDI PSE QB Current Mon 43797
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
--------- ------------ ------
No Alerts to report this report period.
____________________________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 17 September 1999
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ---- -----------------
43724 254/12:03 (11 Sept 99) from 00.0 to 20.6 msec/day
43751 256/07:08 (13 Sept 99) from 20.6 to 00.0 msec/day
43794 259/04:47 (16 Sept 99) from 00.0 to 20.6 msec/day
43814 260/12:06 (17 Sept 99) from 20.6 to 00.0 msec/day
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 17 September 1999
10 Sept 1999 - GMT Day 253 Beta = 47.3 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ---------- -------- ---------- ---- ------ ------
1 -112.0/-134.4 -3.05/-4.31 +0.0/+0.0 16.5 0.000 0.0
2 +39.2/-11.2 +5.79/+4.19 +27.2/-16.0 24.5 1.024 21.0
3 +44.8/-33.6 +4.83/+3.23 +26.8/-16.0 24.3 1.024 20.8
17 Sept 1999 - GMT Day 260 Beta = 27.4 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ---------- -------- ---------- ----- ------- ------
1 -112.0/-134.4 -2.73/-4.31 +0.0/+0.0 16.5 0.000 0.0
2 +33.6/-22.4 +6.44/+5.15 +30.4/-15.2 24.5 1.022 22.0
3 +67.2/-16.8 +5.15/+3.87 +29.6/-16.4 24.5 1.025 21.8
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- ---------- ----------
no changes V/T 5
APPENDIX E
UARS ANOMALY REPORTS
Week ending 17 September 1999
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
99-055 FOT 43768 257/1029 N/A TAC Crashed
99-056 FOT 43778 258/0336 N/A TAC 8 Crashed
99-057 FOT 43809 260/0439 21348 PTP FAILURE, REBOOTED
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
None this report period
ANOMALY INVESTIGATION REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
None this report period
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
99-058 FOT see list Sept 99 14170 Generic Late Acq.
99-055 FOT 43768 257/1029 N/A TAC Crashed
99-056 FOT 43778 258/0336 N/A TAC 8 Crashed
99-057 FOT 43809 260/0439 21348 PTP FAILURE, REBOOTED
SEVNTFW99/38